An electronic transition occurs between two neighboring energy levels and emits radiation with a wavelength of 432 nm. Determine n for the transition. Assume L = 1.2 nm


1
Expert's answer
2021-02-01T03:59:56-0500

Energy expression for particle in a box

En1 = (n12 h2) \ 8mL2

En2 = (n22h2) \ 8mL2

So now emission occoures so n1< n2

Electron jumps from n2 to n1

Energy difference = h2\8mL2 (n22 - n12)

m = 9.11 × 10-31 kg

L = 1.2 nm = 1.2 × 10-9 m

h = 6.626 × 10 34 J. S

Energy difference is = hv = hc\"\lambda"

"\lambda" = 432 nm = 432 × 10-9 m

C = 3 × 108 m s-1

And n2 - n1 = 1




Due to very complex calculation the image is uploaded you can follow very easily and carefully